OpenStudy (anonymous):

Can someone please help me with these?? 3|x+5|<21 2|x-5|<32 Answer choices are -1121 x<-12 or x>2 -12

OpenStudy (anonymous):

\[3|x+5|<21\] step one: divide by 3 what do you get?

OpenStudy (anonymous):

\[3\left[ x+5 \right]<21,\left| x+5 \right|<\frac{ 21 }{ 3 },\left| x+5 \right|<7\]

OpenStudy (anonymous):

\[\left| x \right|<a,means -a<x<a\]

OpenStudy (anonymous):

I still don't get it Can u please do an example?

OpenStudy (anonymous):

\[\left| x-2 \right|<6,-6<x-2<6\]

OpenStudy (anonymous):

you followed.

OpenStudy (anonymous):

I'm lost sorry!

OpenStudy (anonymous):

\[|x-2|<3\iff -3<x-2<3\] and \(2\) all the way across and get \[-1<x<5\]

OpenStudy (anonymous):

\[\left| x \right|<6,-6<x<6\] \[\left| x-2 \right|<6 ,-6<x-2<6, add 2 \to each inequality,-6+2<x-2+2<6+2\] -4<x<8

OpenStudy (anonymous):

i have given two examples.

OpenStudy (anonymous):

Ok thnx

OpenStudy (anonymous):

yw